OpenStudy (anonymous):

what equation represents the axis of symmetry of the funtion y=-2x^2+4x-6

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

the function y=-2x^2+4x-6 has x=1 as its line of symmetry because (let y=f(x)); solve f(x+k)=f(x-k); you will get x=1..... I hope you understand mate.